# Joint Scheduling of DER under Demand Charges: Structure and Approximation

### v1

@@ V-D Simulation Results
As an example, Fig. 4 showed the action sequence for the baseline configuration for HouseZero over one test day. All strategies followed a similar behavior which discharged the battery at the beginning and end of the day. At the same time, demand was reduced over the evening time when there were no solar generations. Both threshold and RL were more aggressive in battery charging actions when lacking generation predictions. The optimal battery actions indicated a much slower charging behavior over the daytime with precise knowledge of solar radiation, creating the performance gap mentioned above. It was also observed that the LSPS algorithm may generate aggressive demand reduction actions due to the lack of predictions. Such behavior negatively impacted the utility function and the overall objective, causing gaps in the optimal solutions.
@@ VI Conclusion
Through simulations with real data from two rich datasets, we demonstrated the performance of the LSPS algorithm in achieving a small solution gap compared to baseline approaches, highlighting its effectiveness compared to relatively more computationally prohibitive methods such as model-predictive control and RL.
The work presented here has several limitations that open avenues for future research. Firstly, the RL algorithm used in our simulations is relatively simple. Designing a more sophisticated algorithm that reliably converges to the theoretical upper bound remains an important direction for future work. Second, the proposed algorithm assumes full knowledge of the utility’s functional form $U(\cdot)$, which may be difficult to obtain in practice. This limitation can be addressed using machine learning or RL techniques that infer the utility function from observed behavior [30].

### v2

@@ V-D Simulation Results
As an example, Fig. 4 shows the action sequence for the baseline configuration of HouseZero over one test day. All strategies followed a similar behavior, which discharged the battery at the beginning and end of the day. At the same time, demand was reduced in the evening when there was no solar generation. Both RATP and RL were more aggressive in battery charging actions when lacking generation predictions. The optimal battery actions indicated a much slower charging behavior over the daytime with precise knowledge of solar generation, creating the performance gap mentioned above. It was also observed that the LSPS algorithm may generate aggressive demand reduction actions due to the lack of predictions. Such behavior negatively impacted the utility function and the overall objective, causing larger solution gaps.
@@ VI Conclusion
Through simulations with real data from two rich and distinct datasets, we demonstrated the performance of the LSPS algorithm in achieving a small solution gap compared to baseline approaches, highlighting its effectiveness compared to relatively more computationally prohibitive methods such as model-predictive control and RL.
The work presented here has several limitations that open avenues for future research. Firstly, the RL algorithm used in our simulations is relatively simple. Designing a more sophisticated algorithm that reliably converges to the theoretical upper bound remains an important direction for future work. Second, the LSPS algorithm assumes full knowledge of the utility’s functional form $U(\cdot)$, which may be difficult to obtain in practice. This limitation can be addressed using machine learning or RL techniques that infer the utility function from observed behavior [35]. Lastly, it is worthwhile to investigate whether sufficient conditions can be established under which the LSPS algorithm remains optimal, even when projections onto the feasible solution set are involved.